Given P(A) = 0.82, P(B) = 0.87, and the fact that events A and B are independent, What is P(A and B)?
Answer by ikleyn(53751)

.

Since events A and B are independent, it means, BY THE DEFINITION, that


the probability of the event  (A and B) is the product  P(A)  and  P(B)

    P(A and B) = P(A)*P(B).


Hence, all you need to do is to multiply the two given numbers.


ANSWER.  P(A and B) = P(A)*P(B) = 0.82*0.87.   Use your calculator.

Solved, with full explanations.